### v2

@@ Abstract
The Arnoldi process provides an efficient framework for approximating functions of a matrix applied to a vector, i.e., of the form $f(M)\bm{b}$, by repeated matrix-vector multiplications. In this paper, we derive an a priori error estimate for approximating the action of a matrix square root using the Arnoldi process, where the integral representation of the error is reformulated in terms of the error for solving the linear system $M\bm{x}=\bm{b}$. The results extend the error analysis of the Lanczos method for Hermitian matrices in [Chen et al., SIAM J. Matrix Anal. Appl., 2022] to non-Hermitian cases. Furthermore, to make the method applicable to large-scale problems, we assume that the matrices are preprocessed utilizing data-sparse approximations preserving positive definiteness, and then establish a refined error bound in this setting. The numerical results on matrices with different structures demonstrate that our theoretical analysis yields a reliable upper bound. Finally, simulations on large-scale matrices arising in particulate suspensions validate the effectiveness and practicality of the approach.
@@ 1 Introduction
Given a matrix $M\in\mathbb{C}^{n\times n}$, its square roots are the solutions of $X^{2}=M.$ If $M$ is positive definite, then $M$ admits a unique positive definite square root [19], denoted as $M^{1/2}$. Matrix square roots frequently arise in stochastic simulation [12], state estimation [2], image reconstruction [3], and Bayesian optimization [31, 18]. Hence, the computation of a matrix square root is a fundamental problem in numerical linear algebra.
@@
There are a list of approaches in the literature for computing the square root of a matrix. Cholesky decomposition and eigenvalue decomposition provide useful tools for symmetric positive definite (SPD) matrices [35]. For general matrices, the Schur decomposition transforms the matrix into a triangular form, where the square root can be computed explicitly [19]. Additionally, Newton-like iterations [19] provide stable and efficient implementation for approximating the matrix square root. Unfortunately, these classical methods all suffer from at least $\mathcal{O}(n^{3})$ complexity for large-scale matrices [19]. On the other hand, even when $M$ is sparse or structured, its square root is typically dense, making direct computation and storage costly or even infeasible for large-scale problems. Consequently, rather than computing the square root of a matrix explicitly, the main interest lies in computing the action of a matrix square root on a vector. This is also the case for general matrix functions. Krylov subspace methods such as the Arnoldi iteration, therefore provide a powerful framework, as they only rely on repeated matrix-vector multiplications to construct the Krylov subspace. For general dense matrices, the overall computational cost of the Arnoldi process is about $\mathcal{O}(kn^{2})$, where $k$ is the number of Arnoldi iterations. When the matrix is Hermitian, the Arnoldi process simplifies to the Lanczos algorithm, characterized by a three-term recurrence. Strategies such as restarting [13, 15, 20], truncated orthogonalization [30, 9], and randomized sketching [30, 9, 29, 16] have been explored to reduce memory and orthogonalization cost.
@@
Recently, there has been growing interest in the theoretical analysis of Krylov methods for approximating the action of matrix functions. As shown in [6], the error of the Lanczos algorithm after $k$ iterations for approximating $f(M)\mathbf{b}$ is bounded by the error of the best $k$-degree polynomial that uniformly approximates $f(x)$ over the spectral interval $[\lambda_{\min}(M),\lambda_{\max}(M)]$, where $\lambda_{\min}(M)$ and $\lambda_{\max}(M)$ denote the smallest and largest eigenvalues of the matrix $M$, respectively. Such a classical bound has also been studied within the setting of finite precision arithmetic [28]. In [11], a posteriori error estimates for polynomial Krylov methods are developed and applied to a broad class of matrix functions with integral representations. Based on the Cauchy integral formula, Chen et al. [6] establish a priori and a posteriori error bounds for the Lanczos method by reformulating the approximation error in terms of the error of solving a shifted linear system. In addition, the authors in [1] explore the near optimality of the Lanczos methods for rational functions with real poles lying outside the spectral interval of $M$. Their results are established under the assumption that the matrices are Hermitian, while in many practical applications, the matrices of interest are non-Hermitian. This motivates us to extend their results to the non-Hermitian setting.
@@
In this work, we present an a priori error bound of the Arnoldi process specifically for evaluating a matrix square root, extending the results in [6] to the case of positive-definite non-Hermitian matrices. The main idea is to utilize the integral representation of the error in the iterations of the Arnoldi process, and then relate it to the error for solving the linear system $M\bm{x}=\bm{b}$. Compared to Hermitian matrices, non-Hermitian matrices exhibit complex spectral behavior: they are not necessarily diagonalizable, and even when they are, the eigenvectors are typically not orthogonal and the eigenvalues may be complex. To address these difficulties, we employ two techniques: the first is applying the Cauchy integral formula to estimate the norms of matrix functions for general matrices that may not be diagonalizable; the second is relating the modulus of a function in the complex plane to the real-valued function of the modulus, reducing the analysis to the positive real line. Moreover, for large-scale problems, we derive an error bound for approximating the matrix square root based on a data-sparse representation of the matrix. In numerical experiments, we investigate matrices with different structures to verify the theoretical results. Furthermore, we validate the effectiveness of our approach on large-scale mobility matrices arising from particulate suspension simulations. To improve computational efficiency and reduce storage cost, the matrices are preprocessed through adaptive cross approximation and stored in the form of hierarchical matrices.
@@
The remainder of this paper is arranged as follows. In Section 2, we introduce the Arnoldi process for matrix functions. Error estimation of the Arnoldi process for computing a square root is provided in Section 3. Section 4 presents the implementation and reports our numerical results. Finally, concluding remarks are made in Section 5.
@@ 2 Matrix functions and Arnoldi iterations
where $Z$ is a nonsingular matrix, and $J=\operatorname{diag}\left(J_{1},J_{2},\ldots,\right.\left.J_{p}\right)$ with $J_{k}\in\mathbb{R}^{m_{k}}$, $\sum_{i=1}^{p}m_{k}=n$ being a Jordan block associated with the eigenvalue $\lambda_{k}$, $1\leq k\leq p$. Then, the matrix function $f(M)$ is defined as
@@
where $\mathcal{C}$ is a closed contour enclosing the spectrum of $M$. The two definitions are equivalent when $f$ is analytic [19].
@@
(6) $MQ_{k}=Q_{k}H_{k}+h_{k+1,k}\bm{q}_{k+1}\bm{e}_{k}^{T},$
Note that when $f(x)=1/x$, the Arnoldi approximation reduces to the Full Orthogonalization Method (FOM) [33] for solving the linear system $M\bm{x}=\bm{b}$, with initial guess $\bm{x}_{0}=\bm{0}$. For the rest of this paper, whenever FOM is referred, we always assume the initial guess is zero.
@@ 3 Error bound
In this section, we study an a priori error bound of the Arnoldi process for approximating general matrix functions and then focus on the analysis of the matrix square root.
@@ 3.1 Integral representation of the error
Assuming $M$ and $M-zI$ are invertible for some $z\in\mathbb{C}$, we obtain
$\bm{\xi}_{z}^{k}=\det(h_{z}(H_{k}))h_{z}(M)\bm{\xi}_{0}^{k}\quad\text{and}\quad\bm{r}_{z}^{k}=\det(h_{z}(H_{k}))\bm{r}_{0}^{k},$
where the function $h_{z}(x)$ is defined as $h_{z}(x):=(x-z)^{-1}x$.
@@
The proof follows a similar argument to that of Corollary 2.4 in [6].
@@
Given a nonsingular matrix $M\in\mathbb{C}^{n\times n}$ and a vector $\bm{b}\in\mathbb{C}^{n}$, let function $f$ be analytic on and inside a closed contour $\mathcal{C}$ enclosing the spectrum of $M$ and $H_{k}$, where $H_{k}$ is an upper Hessenberg matrix generated from the Arnoldi process for $M$ and $\bm{b}$. Then, the error of the Arnoldi iteration for $f(M)b$ can be expressed as
(8) $f(M)\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})=\left(-\frac{1}{2\pi i}\int_{\mathcal{C}}f(z)\operatorname{det}\left(h_{z}\left(H_{k}\right)\right)h_{z}(M)\,\mathrm{d}z\right)\bm{\xi}_{0}^{k}.$
@@
$=-\frac{1}{2\pi i}\int_{\mathcal{C}}f(z)\det(h_{z}(H_{k}))h_{z}(M)\bm{\xi}_{0}^{k}\,\mathrm{d}z.$
@@
Theorem 3.5 shows that for any nonsingular $M$, the error of the Arnoldi iteration for computing $f(M)\bm{b}$ is reformulated as the error in solving the linear system $M\bm{x}=\bm{b}$. This further simplifies the contour integral. For a more general setting, the error admits a bound related to the error of solving shifted linear systems of the form $(M-wI)\bm{x}=\bm{b}$, where $w$ lies outside the spectrum of both $H_{k}$ and $M$; see Corollary 2.5 in [6] for details.
@@ 3.2 Matrix square root
We now turn to deriving an a priori error estimate of the Arnoldi process for approximating a matrix square root, that is $f(x)=x^{1/2}$. To ensure the existence of a unique positive definite square root, we assume that $M$ is positive definite. Here, the matrices of interest are non-Hermitian. For such matrices, the eigenvalues may lie off the real axis leading to a contour integral in the complex plane for the error of the Arnoldi process, thereby introducing challenges. We derive the error bound starting from Theorem 3.5, and organize the proof into the following steps:
Step 1: Select an appropriate contour in the complex plane, and derive an initial error bound for the 2-norm of the Arnoldi approximation.
Step 2: Bound the modulus of the complex function $h_{z}(w)$ using a real-valued function of the modulus, and then estimate $\|h_{z}(M)\|$ with a new contour integral.
Step 3: Apply the properties of the Beta and Gamma functions to obtain an error bound in terms of the residual $\bm{r}_{0}^{k}$ for solving $Mx=b$.
Step 4: Conclude the proof by deriving a upper bound for $\bm{r}_{0}^{k}$.
From this, we establish the following theorem.
@@
Given a positive-definite matrix $M\in\mathbb{C}^{n\times n}$ and a vector $\bm{b}\in\mathbb{C}^{n}$, the relative error of the Arnoldi process for approximating the matrix square root is bounded by
$\frac{\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\bm{b})\|}{\|M^{1/2}\bm{b}\|}\leq\frac{2\sqrt{2}(\kappa(M))^{5/2}}{(k-1/2)^{3/4}}\frac{\|\bm{r}^{k}_{0}\|}{\|\bm{b}\|}\leq\frac{4\sqrt{2}(\kappa(M))^{7/2}}{(k-1/2)^{3/4}},$
where $\kappa(M)$ denotes the condition number of $M$, defined as $\kappa(M)=\frac{\sigma_{\max}(M)}{\sigma_{\min}(M)}$, and $\sigma_{\max}(M)$ and $\sigma_{\min}(M)$ denote the largest and smallest singular values of $M$, respectively
@@
Step 1: Choose an Appropriate Contour. Define $\mathcal{C}$ as a keyhole contour, consisting of a large circular arc $\mathcal{C}_{R}$ of radius $R$ centered at the origin, a small indented circular arc $\mathcal{C}_{\epsilon}$ of radius $\epsilon$ around the origin, and two horizontal line segments $\mathcal{C}^{+}_{\delta}$ and $\mathcal{C}^{-}_{\delta}$, which run at a small vertical distance $0<\delta<\epsilon$ above and below the branch cut, respectively (see Fig. 1 for details).
@@
$\left\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})\right\|\leq$ $\left(\frac{1}{2\pi}\int_{\mathcal{C}_{R}}|z^{1/2}|\cdot\left|\operatorname{det}\left(h_{z}\left(H_{k}\right)\right)\right|\cdot\left\|h_{z}(M)\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|$
$-\left(\frac{1}{2\pi}\int_{\mathcal{C}_{\epsilon}}|z^{1/2}|\cdot\left|\operatorname{det}\left(h_{z}\left(H_{k}\right)\right)\right|\cdot\left\|h_{z}(M)\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|$
$+\left(\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{+}}|z^{1/2}|\cdot\left|\operatorname{det}\left(h_{z}\left(H_{k}\right)\right)\right|\cdot\left\|h_{z}(M)\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|$
(9) $-\left(\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{-}}|z^{1/2}|\cdot\left|\operatorname{det}\left(h_{z}\left(H_{k}\right)\right)\right|\cdot\left\|h_{z}(M)\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|.$
Let $\mathcal{I}_{1},\ldots,\mathcal{I}_{k}\subset\mathbb{C}$ be subsets such that the eigenvalues of $H_{k}$ are contained in $\mathcal{I}_{i}$ for $i=1,\ldots,k$. Define the norm $\|h\|_{\mathcal{I}}:=\sup_{x\in\mathcal{I}}|h(x)|$ for a function $h(x)$ on $\mathcal{I}\subset\mathbb{C}$, where $|h(x)|$ denotes the modulus of $h(x)$. Note that the Jordan block definition of matrix functions Eq. 5 implies the eigenvalues of $f(M)$ are exactly $f(\lambda_{i})$, where $\lambda_{i}$ are the eigenvalues of $M$. Thus,
$\left|\operatorname{det}\left(h_{z}\left(H_{k}\right)\right)\right|=\left|\prod_{i=1}^{k}\lambda_{i}\left(h_{z}\left(H_{k}\right)\right)\right|=\left|\prod_{i=1}^{k}h_{z}\left(\lambda_{i}\left(H_{k}\right)\right)\right|\leq\prod_{i=1}^{k}\left\|h_{z}\right\|_{\mathcal{I}_{i}}.$
Plugging this into (3.9) yields
$\left\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})\right\|\leq$ $\left(\frac{1}{2\pi}\int_{\mathcal{C}_{R}}|z^{1/2}|\cdot\prod_{i=1}^{k}\left\|h_{z}\right\|_{\mathcal{I}_{i}}\cdot\left\|h_{z}(M)\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|$
$-\left(\frac{1}{2\pi}\int_{\mathcal{C}_{\epsilon}}|z^{1/2}|\cdot\prod_{i=1}^{k}\left\|h_{z}\right\|_{\mathcal{I}_{i}}\cdot\left\|h_{z}(M)\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|$
$+\left(\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{+}}|z^{1/2}|\cdot\prod_{i=1}^{k}\left\|h_{z}\right\|_{\mathcal{I}_{i}}\cdot\left\|h_{z}(M)\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|$
(10) $-\left(\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{-}}|z^{1/2}|\cdot\prod_{i=1}^{k}\left\|h_{z}\right\|_{\mathcal{I}_{i}}\cdot\left\|h_{z}(M)\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|.$
As the outer radius $R\to\infty$, the integral over the large circular arc vanishes for $k\geq 1$ since $|z^{1/2}|=O\left(R^{1/2}\right)$, $\left\|h_{z}\right\|_{\mathcal{I}_{i}}=O\left(R^{-1}\right)$, and $\|h_{z}(M)\|=O(R^{-1})$. Similarly, as the inner radius $\epsilon\to 0$, the integral goes to zero. Thus, as $\epsilon\to 0$ and $R\to\infty$, the contributions to the integral come from $\mathcal{C}^{+}_{\delta}$ and $\mathcal{C}^{-}_{\delta}$. Also, as $\epsilon\to 0$, $\delta$ approaches 0. Then, we have
@@
$\leq$ $\lim_{\begin{subarray}{c}\epsilon\to 0\\ R\to\infty\end{subarray}}\left(\frac{1}{2\pi}\int_{-R}^{-\epsilon}\left|(x+\delta i)^{1/2}\right|\cdot\left(\prod_{i=1}^{k}\left\|h_{x+\delta i}\right\|_{\mathcal{I}_{i}}\right)\cdot\left\|h_{x+\delta i}(H_{k})\right\|\mathrm{~{}d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
@@
$-\left(\frac{1}{2\pi}\int^{-R}_{-\epsilon}\left|(x-\delta i)^{1/2}\right|\cdot\left(\prod_{i=1}^{k}\left\|h_{x-\delta i}\right\|_{\mathcal{I}_{i}}\right)\cdot\left\|h_{x-\delta i}(H_{k})\right\|\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
$=$ $\left(\frac{1}{\pi}\int_{0}^{\infty}x^{1/2}\cdot\left(\prod_{i=1}^{k}\left\|h_{-x}\right\|_{\mathcal{I}_{i}}\right)\cdot\left\|h_{-x}(M)\right\|\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|.$
Step 2: Bound $\left\|h_{-x}\right\|_{\mathcal{I}_{i}}$ and $\left\|h_{-x}(M)\right\|$. Consider any $w\in\mathbb{C}$ and a fixed $x\in\mathbb{R}$ such that $x>0$, then,
$|w+x|=\sqrt{(\operatorname{Re}(w)+x)^{2}+\operatorname{Im}(w)^{2}}\quad\text{and}\quad\sqrt{|w|^{2}+x^{2}}=\sqrt{\operatorname{Re}(w)^{2}+\operatorname{Im}(w)^{2}+x^{2}}.$
A straightforward calculation then yields
$|h_{-x}(w)|=\frac{|w|}{|w+x|}\leq\frac{|w|}{\sqrt{|w|^{2}+x^{2}}}:=g_{x}(w).$
The function $g_{x}(w)$ is increasing with respect to $|w|$. Noting that $H_{k}=Q_{k}^{*}MQ_{k}$, we apply the interlacing inequalities for singular values in Theorem 2 of [34] to obtain:
$\sigma_{\max}(H_{k})\leq\sigma_{\max}(M)\quad\text{and}\quad\sigma_{\min}(H_{k})\geq\sigma_{\min}(M).$
Furthermore, the modulus of the eigenvalues of a matrix is bounded above by its largest singular value, and the modulus of the smallest eigenvalue is bounded below by its smallest singular value. Letting $\mathcal{S}:=[\sigma_{\min}(M),\sigma_{\max}(M)]$, the problem now turns into analyzing the behavior of $g_{x}$ within $\mathcal{S}$,
(11) $\left\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})\right\|\leq\left(\frac{1}{\pi}\int_{0}^{\infty}x^{1/2}\cdot\left(\prod_{i=1}^{k}\left\|g_{x}\right\|_{\mathcal{S}}\right)\cdot\left\|h_{-x}(M)\right\|\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|.$
We next give the bound for $\left\|h_{-x}(M)\right\|$. For a fixed real number $x>0$, we consider a new contour $\mathcal{C}$ as in Fig. 1, but with $R=2\sigma_{\max}(M)$ and $\delta=\frac{\epsilon C_{M}}{\sqrt{C_{M}^{2}+4}}$, where $C_{M}=\frac{\lambda_{\min}\left(\frac{M+M^{*}}{2}\right)}{\left|\lambda_{\min}\left(i\left(\frac{M-M^{*}}{2}\right)\right)\right|}$. Using the Cauchy integral formula, it follows that
$h_{-x}(M)=-\frac{1}{2\pi i}\int_{\mathcal{C}}h_{-x}(s)(M-sI)^{-1}\,\mathrm{d}s.$
Then, we have:
$\left\|h_{-x}(M)\right\|$
$\leq$ $\frac{1}{2\pi}\int_{\mathcal{C}_{R}}|h_{-x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|-\frac{1}{2\pi}\int_{\mathcal{C}_{\epsilon}}|h_{-x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|$

### v3

@@ Abstract
The Arnoldi process provides an efficient framework for approximating functions of a matrix applied to a vector, i.e., of the form $f(M)\bm{b}$, by repeated matrix-vector multiplications. In this paper, we derive error estimates for approximating the action of a matrix square root using the Arnoldi process, where the integral representation of the error is reformulated in terms of the error for solving the linear system $M\bm{x}=\bm{b}$. The results extend the error analysis of the Lanczos method for Hermitian matrices in [Chen et al., SIAM J. Matrix Anal. Appl., 2022] to non-Hermitian cases and provide an improved bound for the Hermitian case. Furthermore, in practical settings, the matrix may only be available via approximate or structured representations. Motivated by this, we extend the analysis and establish a generalized error bound for perturbed matrices. The numerical results on matrices with different structures demonstrate that our theoretical analysis yields a reliable upper bound. Finally, simulations on large-scale matrices arising in particulate suspensions, represented in hierarchical matrix form, validate the effectiveness and practicality of the approach.
@@ 1 Introduction
Given a matrix $M\in\mathbb{C}^{n\times n}$, its square roots are the solutions of $X^{2}=M.$ If $M$ is positive definite, then $M$ admits a unique positive definite square root [higham2008functions], denoted as $M^{1/2}$. Matrix square roots frequently arise in stochastic simulation [ColloidsReview_COCIS2019], state estimation [asl2019adaptive], image reconstruction [bardsley2012mcmc], and Bayesian optimization [pleiss2020fast, hernandez2014predictive]. Hence, the computation of a matrix square root is a fundamental problem in numerical linear algebra.
@@
There are a list of approaches in the literature for computing the square root of a matrix. Cholesky decomposition and eigenvalue decomposition provide useful tools for symmetric positive definite (SPD) matrices [watkins2004fundamentals]. For general matrices, the Schur decomposition transforms the matrix into a triangular form, where the square root can be computed explicitly [higham2008functions]. Additionally, Newton-like iterations [higham2008functions] provide stable and efficient implementation for approximating the matrix square root. However, these classical methods all suffer from at least $\mathcal{O}(n^{3})$ complexity for large-scale matrices [higham2008functions]. Even when $M$ is sparse or structured, its square root is typically dense, making direct computation and storage costly or even infeasible for large-scale problems. Consequently, rather than computing the square root of a matrix explicitly, one can instead compute the action of a matrix square root on a vector. This is also the case for general matrix functions. Krylov subspace methods such as the Arnoldi iteration, therefore provide a powerful framework, as they only rely on repeated matrix-vector multiplications to construct the Krylov subspace. For general dense matrices, the overall computational cost of the Arnoldi process is about $\mathcal{O}(kn^{2})$, where $k$ is the number of Arnoldi iterations. When the matrix is Hermitian, the Arnoldi process simplifies to the Lanczos algorithm [chen2024lanczos], characterized by a three-term recurrence.
@@
There has been growing interest in the study and discussion of Krylov methods for approximating the action of matrix functions. An early study by Saad [saad1992analysis] related the error of Krylov subspace approximations for $f(M)\mathbf{b}$ to polynomial approximation. If $M$ is Hermitian, then one obtains the near-optimality property of Lanczos, that is, the error after $k$ iterations is bounded by that of the best polynomial of degree less than $k$ uniformly approximating $f(x)$ over the spectrum of $M$; see, e.g. [musco2018stability, chen2022error]. In [diele2009error], the error of polynomial Krylov approximations for matrix functions is represented in integral forms, and a posteriori error estimates for exponential-like and trigonometric matrix functions are provided. In [eiermann2006restarted, ilic2010restarted], the authors expressed the errors of Krylov subspace approximations in terms of divided differences, providing insights into the design of restart schemes. For general analytic functions $f$ admitting an integral representation, [frommer2014efficient] provided an integral representation for the error of the Arnoldi approximation to $f(M)\bm{b}$, and proposed a restarting strategy based on evaluating this error function via numerical quadrature. Based on [frommer2014efficient, frommer2014convergence], Frommer and Schweitzer [frommer2016error] derived error bounds and estimates of the Krylov subspace approximations $f(M)\bm{b}$ for Stieltjes functions. Using the Cauchy integral formula, Chen et al. [chen2022error] established a priori and a posteriori error bounds for the Lanczos method by reformulating the approximation error in terms of the error of solving a shifted linear system. In addition, Amsel et al. [amsel2023near] explored the near optimality of the Lanczos methods for rational functions with real poles lying outside the spectral interval of $M$. Recently, [schweitzer2025near] proved that, for Stieltjes and related functions, the Lanczos method exhibits, what they call, near-instance optimality. There are also several works that studied Krylov approximations in the context of rational approximations [frommer2009error, frommer2008stopping, frommer20132] and matrix exponentials [beckermann2009error, stewart1996error, hochbruck1997krylov, druskin1998using]. Most of these results have been established under the assumption that the matrices are Hermitian, while in many practical applications, the matrices of interest are non-Hermitian, which motivates us to extend the results to the non-Hermitian setting.
@@
In this work, we present error bounds of the Arnoldi process specifically for evaluating a matrix square root, extending the results in [chen2022error] to the case of positive-definite non-Hermitian matrices. In this context, we say that $M$ is positive-definite if $\operatorname{Re}(\bm{x}^{*}M\bm{x})>0$ for any unit vector $\bm{x}\in\mathbb{C}^{n}$. The main idea is to utilize the integral representation of the error in the iterations of the Arnoldi process, and then relate it to the error for solving the linear system $M\bm{x}=\bm{b}$. Compared to Hermitian matrices, non-Hermitian matrices exhibit complex spectral behavior: they are not necessarily diagonalizable, and even when they are, the eigenvectors are typically not orthogonal and the eigenvalues may be complex. To address these difficulties, we relate the modulus of a function in the complex plane to the real-valued function of the modulus, reducing the analysis to the positive real line. This also allows us to derive a sharper bound for the Hermitian case by using the average of eigenvalues, which improves on the results in [chen2022error, amsel2023near]. Additionally, for large-scale problems, we derive an error bound for approximating the matrix square root based on a data-sparse representation of the matrix. In numerical experiments, we investigate matrices with different structures to verify the theoretical results. Finally, we validate the effectiveness of our approach on large-scale mobility matrices arising from particulate suspension simulations. To improve computational efficiency and reduce storage cost, the matrices are preprocessed through adaptive cross approximation and stored in the form of hierarchical matrices.
@@
The remainder of this paper is organized as follows. In Section 2, we introduce the Arnoldi process for matrix functions. The error estimation of the Arnoldi process for computing a square root is provided in Section 3. In Section 4, numerical examples are presented to validate the theoretical results discussed in this paper. Finally, concluding remarks are provided in Section 5.
@@ 2 Matrix functions and Arnoldi iterations
where $Z$ is a nonsingular matrix, and $J=\operatorname{diag}\left(J_{1},J_{2},\ldots,\right.\left.J_{p}\right)$ with $J_{k}\in\mathbb{R}^{m_{k}\times m_{k}}$, $\sum_{i=1}^{p}m_{k}=n$, being a Jordan block associated with the eigenvalue $\lambda_{k}$, $1\leq k\leq p$. Then, the matrix function $f(M)$ is defined as
@@
where $\mathcal{C}$ is a closed contour enclosing the spectrum of $M$. The two definitions are equivalent when $f$ is analytic [higham2008functions].
(6) $MQ_{k}=Q_{k}H_{k}+h_{k+1,k}\bm{q}_{k+1}\bm{e}_{k}^{T},$
@@
Note that when $f(x)=1/x$, the Arnoldi approximation reduces to the Full Orthogonalization Method (FOM) [saad2003iterative] for solving the linear system $M\bm{x}=\bm{b}$, with initial guess $\bm{x}_{0}=\bm{0}$. For the rest of this paper, whenever FOM is referred, we always assume the initial guess is zero.
@@ 3 Error bound
In this section, we study error bounds of the Arnoldi process for approximating general matrix functions and then focus on the analysis of the matrix square root.
@@ 3.1 Integral representation of the error
Following a similar argument to that in [chen2022error], we have the following results.
@@
Assuming $M$ and MzIM-zI are invertible for some $z\in\mathbb{C}$, we obtain
$\bm{\xi}_{z}^{k}=\frac{\det(H_{k})}{\det(H_{k}-zI)}(M-zI)^{-1}M\bm{\xi}_{0}^{k}\quad\text{and}\quad\bm{r}_{z}^{k}=\frac{\det(H_{k})}{\det(H_{k}-zI)}\bm{r}_{0}^{k}.$
@@
The proof follows a similar argument to that of Corollary 2.4 in [chen2022error].
@@
Given a nonsingular matrix $M\in\mathbb{C}^{n\times n}$ and a vector $\bm{b}\in\mathbb{C}^{n}$, let function $f$ be analytic on and inside a closed contour $\mathcal{C}$ enclosing the spectrum of $M$ and $H_{k}$, where $H_{k}$ is an upper Hessenberg matrix generated from the Arnoldi process for $M$ and $\bm{b}$. Then, the error of the Arnoldi iteration for $f(M)\bm{b}$ can be expressed as
(8) $f(M)\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})=\left(-\frac{1}{2\pi i}\int_{\mathcal{C}}f(z)\frac{\det(H_{k})}{\det(H_{k}-zI)}(M-zI)^{-1}M\,\mathrm{d}z\right)\bm{\xi}_{0}^{k}.$
@@
$=-\frac{1}{2\pi i}\int_{\mathcal{C}}f(z)\frac{\det(H_{k})}{\det(H_{k}-zI)}(M-zI)^{-1}M\bm{\xi}_{0}^{k}\,\mathrm{d}z.$
@@
Here we assume that the function $f$ is analytic on and inside a closed contour $\mathcal{C}$ enclosing the spectrum of $M$ and $H_{k}$. However, the theorem holds under the sufficient condition that $\mathcal{C}$ encloses the field of values of $M$.
Similarly to the results in [frommer2009error, frommer2014efficient], Theorem 3.5 shows that for any nonsingular $M$, the error of the Arnoldi iteration for computing $f(M)\bm{b}$ is reformulated as the error in solving the linear system $M\bm{x}=\bm{b}$. This further simplifies the contour integral. For a more general setting, the error admits a bound related to the error of solving shifted linear systems of the form $(M-wI)\bm{x}=\bm{b}$, where $w$ lies outside the spectrum of both $H_{k}$ and $M$.
@@ 3.2 Matrix square root
We now turn to deriving error estimates of the Arnoldi process for approximating a matrix square root, that is $f(x)=x^{1/2}$. To ensure the existence of a unique positive definite square root, we assume that $M$ is positive-definite.
@@
Given a positive-definite matrix $M\in\mathbb{C}^{n\times n}$ and a vector $\bm{b}\in\mathbb{C}^{n}$, the error of the Arnoldi process for approximating the matrix square root is bounded by
$\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\bm{b})\|\leq\frac{1}{\pi}\left(\int_{0}^{\infty}x^{1/2}\prod_{i=1}^{k}\left|\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})+x}\right|\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|,$
where $\lambda_{1}(H_{k})\geq\lambda_{2}(H_{k})\geq\cdots\geq\lambda_{k}(H_{k})$ are the eigenvalues of $H_{k}$.
@@
Define $\mathcal{C}$ as a keyhole contour, consisting of a large circular arc $\mathcal{C}_{R}$ of radius $R$ centered at the origin, a small circular arc $\mathcal{C}_{\epsilon}$ of radius $\epsilon$ around the origin, and two horizontal line segments $\mathcal{C}^{+}_{\delta}$ and $\mathcal{C}^{-}_{\delta}$, which run at a small vertical distance $0<\delta<\epsilon$ above and below the branch cut, respectively (see Fig. 1 for details).
@@
Using properties of the determinant, we have
$\left|\frac{\det(H_{k})}{\det(H_{k}-zI)}\right|=\left|\prod_{i=1}^{k}\frac{\lambda_{i}\left(H_{k}\right)}{\lambda_{i}(H_{k}-zI)}\right|=\prod_{i=1}^{k}\left|\frac{\lambda_{i}\left(H_{k}\right)}{\lambda_{i}(H_{k})-z}\right|.$
@@
$\leq$ $\left(\frac{1}{2\pi}\int_{\mathcal{C}_{R}}|z^{1/2}|\prod_{i=1}^{k}\left|\frac{\lambda_{i}\left(H_{k}\right)}{\lambda_{i}(H_{k})-z}\right|\left\|(M-zI)^{-1}M\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|$
@@
$-\left(\frac{1}{2\pi}\int_{\mathcal{C}_{\epsilon}}|z^{1/2}|\prod_{i=1}^{k}\left|\frac{\lambda_{i}\left(H_{k}\right)}{\lambda_{i}(H_{k})-z}\right|\left\|(M-zI)^{-1}M\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|$